show me the steps to solve 3(7-4(-3x-6))=36x+93
Answer by kekegeter (97)

3(-(-4*3x-4*6)+7)=36x+93
3(-(-12x-24)+7)=36x+93
3(12x+24+7)=36x+93
3*12x+3*31=36x+93
...
Just in case you cannot get it, the answer is 36x+93+ = 36x+93